OpenStudy (anonymous):

What number is 56%. Of 48

OpenStudy (anonymous):

27.84

OpenStudy (anonymous):

56*48/100

OpenStudy (anonymous):

@ghass1878 should make it more clear to her >>>> (56*48)/ 100

OpenStudy (anonymous):

ok:)

OpenStudy (cornzyblack):

(56/100) * 48 = 0.56*48= 26.88